### v1

Abstract
Full Waveform Inversion (FWI) is a widely used method in seismic data processing, capable of estimating models that represent the characteristics of the geological layers of the subsurface. Because it works with a massive amount of data, the execution of this method requires much time and computational resources, being restricted to large-scale computer systems such as supercomputers. Techniques such as FWI adapt well to parallel computing and can be parallelized in shared memory systems using the application programming interface (API) OpenMP. The management of parallel tasks can be performed through loop schedulers contained in OpenMP. The dynamic scheduler stands out for distributing predefined fixed-size chunk sizes to idle processing cores at runtime. It can better adapt to FWI, where data processing can be irregular. However, the relationship between the size of the chunk size and the runtime is unknown. Optimization techniques can employ meta-heuristics to explore the parameter search space, avoiding testing all possible solutions. Here, we propose a strategy to use the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA), with Coupled Simulated Annealing (CSA) as its optimization method, to automatically adjust the chunk size for the dynamic scheduling of wave propagation, one of the most expensive steps in FWI. Since testing each candidate chunk size in the complete FWI is unpractical, our approach consists of running a PATSMA where the objective function is the runtime of the first time iteration of the first seismic shot of the first FWI iteration. The resulting chunk size is then employed in all wave propagations involved in an FWI. We conducted tests to measure the runtime of an FWI using the proposed autotuning, varying the problem size and running on different computational environments, such as supercomputers and cloud computing instances. The results show that applying the proposed autotuning in an FWI reduces its runtime by up to 70.46% compared to standard OpenMP schedulers.
1 Introduction
Seismic surveying is one of the leading geophysical techniques. This technique is used to identify subsurface features, with the ultimate goal of obtaining an image of the subsurface. This method is widespread in the exploration of oil and gas. Input data used in seismic surveying are obtained through seismic acquisition. In this stage, seismic shots, typically near the surface, generate seismic waves that propagate through the subsurface and can be reflected. Sensors, such as geophones or hydrophones, can record the energy of such waves as seismic traces. The collection of such traces is called observed data. They are processed to obtain information from the subsurface.
Seismic modeling can be one of the processing steps. It is responsible for simulating the propagation of waves through the subsurface. This simulation generates information about how seismic waves propagate through the subsurface and how they are affected by the geophysical properties of the subsurface, such as the velocity of seismic wave propagation. The simulation is based on geophysical models of the subsurface developed based on the observed data and other geological information. Many aspects, such as density, acoustic pressure, and particle velocity, can be modeled.
The more modeled parameters, the more accurate the seismic modeling is. On the other hand, increasing the number of modeled parameters can make the computation more complex and time-consuming. The reason is that the simulation of seismic wave propagation involves intensive mathematical calculations that require a lot of computational processing power. The information generated from seismic modeling is called calculated data.
The procedure to obtain a velocity model from seismic data is called velocity analysis. Full Waveform Inversion (FWI) (Tarantola, 1984) is one of the most used methods to estimate a velocity model. An FWI iteratively adjusts the velocity model using an optimization strategy. It aims to reduce the residual data, i.e., the difference between the observed and calculated data. FWI computes adjustments to the velocity models from the calculated and residual data.
FWI also presents challenges, such as the need for high-quality data and the computational complexity of the process. Due to its high computational demands, the execution of this algorithm in industry models is restricted to large computing environments where the use of parallel programming techniques is necessary.
OpenMP (Open Multi-Processing ) (Dagum and Menon, 1998), a library for parallel programming in shared memory systems, is a widely used tool in the parallelization of large-scale applications such as the FWI. It allows programmers to write parallel code to distribute the work of an application between multiple execution threads. Those threads are then assigned to cores of a shared memory system. By default, in OpenMP, this load is divided equally. However, in complex problems such as FWI, a common challenge is load imbalance, where some parts of the code execute unevenly regular tasks, making some threads complete their work more quickly than others. The faster threads will become idle while waiting for the other threads to finish. Load imbalance can result in underutilization of the available hardware and, consequently, a reduction in the program’s overall performance.
Here, we propose a method to balance the workload between CPUs on a computing node, avoiding resource underutilization. To do this, we employ the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) (Fernandes et al., 2024) to automatically adjust the chunk size of OpenMP dynamic scheduler. The main contribution of this work is to adapt PATSMA for FWI and evaluate its performance. We compare our proposal with standard OpenMP schedulers, showing that the proposed autotuning reduces the FWI execution time in different computing environments testing for several input sizes.
The rest of this paper is organized as follows: Section 2 presents the characteristics of our target application, FWI, and the autotuning method we use, PATSMA. It also discusses workload distribution through loop scheduling supported by OpenMP. Next, in Section 3, we detail our FWI application and the proposed autotuning approach. We show the proposed method’s results compared with the standard OpenMP schedulers in Section 4. Section 5 presents the related work, and Section 6 concludes this work.
2 Background
This section briefly describes our target application, the FWI (Section 2.1), and the autotuning method we use, PATSMA (Section 2.2). It also presents the standard OpenMP loop schedulers we compare to our approach (Section 2.3).
2.1 FWI
Full waveform inversion (FWI) (Tarantola, 1984) is a seismic data inversion technique that aims to obtain an accurate model of the subsurface geophysical properties. This technique minimizes an objective function that measures the difference between observed and calculated data, i.e., the residual data. The gradient of the objective function indicates the search direction for the FWI optimization problem (Plessix, 2006). The model is updated based on that gradient. Mathematically, FWI can be described as a least squares optimization problem whose objective function is
$\min_{\mathbf{m}\in\mathbb{R}^{n}}\frac{1}{2}\|\mathbf{d-G(m)}\|^{2},$ (1)
[…]
This process is repeated until a stopping criterion is reached. Common stopping conditions are the convergence of the optimization method (the norm of the gradient is less than a predefined threshold) or the maximum number of iterations. The main steps of FWI are:
  1. 1.
    Read observed data;
  2. 2.
    Read an initial model;
  3. 3.
    Compute the solution of the wave equation (calculated data) for the initial model;
  4. 4.
    Compute the objective function (Equation 1);
  5. 5.
    Compute the gradient of the objective function;
  6. 6.
    Update the model based on the gradient of the objective function;
  7. 7.
    Repeat steps 3 to 6 until the stopping criterion is reached.
2.2 PATSMA
The Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) is a software that aims to adjust a parameter of a shared memory application using an optimization method. The application programmer can define both the parameter and the optimization method. Although PATSMA is suitable for different optimization methods, its software package includes Coupled Simulated Annealing (CSA) (Xavier-de Souza et al., 2010).
Coupled Simulated Annealing (CSA) is a global optimization method that can find optimal solutions for problems with multiple local minima. CSA is an extension of the Simulated Annealing (SA) algorithm (Kirkpatrick et al., 1983) that uses multiple SA instances to explore the solution space simultaneously. In CSA, two or more optimizers are coupled, i.e., at each iteration, the optimizers exchange information about the best solution.
The SA algorithm is summarized in two phases: the generation of a new solution and the acceptance of that solution. Some parameters control both process stages, known as the generation and acceptance temperatures. A generation temperature function obtains new candidate solutions. When a candidate solution has a lower cost than the previous solution, it is accepted with a probability of $100\%$. Otherwise, that solution is accepted as the new solution with a probability calculated using an acceptance temperature function.
CSA increases the search diversity by using multiple SA processes in different subsets of the search space. The exchange of information between the optimizers helps to diversify the search, allowing CSA to find global solutions more efficiently. Additionally, this exchange enables automatic adjustment of the acceptance temperature, making CSA less dependent on the initial value of this parameter.
The optimizers can start from different starting points in the search space. As a result, CSA can explore the solution space more widely than standard SA. At each iteration, each optimizer generates a new solution. If that solution worsens, it can still be selected according to some defined probability. The algorithm becomes less likely to accept a worse solution as the process continues. CSA can be a very effective technique for finding optimal solutions to global optimization problems. However, precisely tuning the CSA parameters is essential to ensure that CSA can generate satisfactory solutions for specific problems. The advantage of CSA over SA is that CSA is less sensitive to the choice of parameters and can produce solutions of better quality even when the parameters are not tuned optimally.
2.3 OpenMP schedulers
OpenMP (Open Multi-Processing) is an Application Programming Interface (API) for parallel programming on multicore systems with shared memory (Board, 2021). It provides tools for programmers to develop parallel codes in C, C++, and Fortran without significant changes to the original code, using its directives to distribute application work across multiple threads.
#pragma omp for is a widely used OpenMP directive. That directive distributes the iterations of a loop between the threads that execute them simultaneously in different cores, allowing the program to execute these loops faster in shared memory systems. schedule is one of the most important clauses of the directive omp for. A clause is an instruction to change the default behavior of a directive. The clause schedule specifies how the loop iterations must be distributed among the threads. It provides three default schedulers: static, dynamic, and guided. The operation of each default scheduler is defined below.
  • static: Chunks of iterations are distributed round-robin among the threads before the loop is executed. The programmer may define the size of the chunks. If the chunk size is not defined, its value is approximately $N_{i}/N_{t}$, where $N_{i}$ is the number of iterations and $N_{t}$ is the number of threads.
  • dynamic: During runtime, chunks of loop iterations are distributed among the threads. When a thread finishes executing its chunk, it requests a new chunk from a global queue. If the programmer does not specify the chunk size, it is $1$.
  • guided: Similar to dynamic, the chunks of iterations are distributed at runtime. However, the chunk size decreases with each new request. This strategy aims to handle load imbalance better. When the programmer specifies the chunk size, no chunk can be smaller than that except the last one. If a chunk size is not defined, it is $1$.
The static scheduler can be advantageous because static distribution guarantees that each thread will execute approximately the same number of iterations. As there is no workload redistribution overhead during execution, that strategy may lead to better performance for homogeneous hardware and iterations workload. However, if the iterations workload or the hardware is heterogeneous, static distribution may result in resource underutilization as some threads may finish before others.
Applications with heterogeneous iterations workload or running in heterogeneous hardware may perform better using the dynamic scheduler. Because such a scheduler distributes the chunks of iterations at runtime, no thread stays idle until all the chunks are distributed. However, the program may lose performance if the block size is not chosen appropriately. If the chunk is too large, it can generate idle time. On the other hand, if the chunk is too small, there may be excessive chunk requests, leading to a significant increase in runtime.
The guided scheduler aims to reduce idleness compared to the static scheduler and management overhead compared to the dynamic scheduler. Because the last chunks are smaller, threads will not idle long. Also, as the first chunks are larger, threads will be busy longer, reducing the number of new chunk requests. The disadvantages of the guided scheduler are that (1) there is an additional overhead to manage the chunk size, and (2) an autotuning method could only control the size of the last block.
Therefore, finding an ideal chunk size is essential to achieving the best possible performance of loop scheduling of an application using OpenMP. That way, it is critical to carefully adjust the chunk size for each application, depending on the number of loop iterations, the workload characteristics, and the computing environment.
Assis et al. (2020) evaluated the performance of the dynamic and static schedulers using a variety of simulations in a reverse time migration (RTM) algorithm (Baysal et al., 1983). The results showed that the dynamic scheduler with chunk size autotuning consistently outperformed the standard static and guided schedulers, with improvements of up to $33\%$ in some cases. Therefore, we used the dynamic scheduler in this work.
3 Autotuning applied to FWI
An algorithm to calculate the solution of the 3D wave equation through finite differences comprises nested loops representing the space domain. The solution of the wave equation is one of the most time-consuming parts of an FWI and is executed several times. Those loops occur in FWI in the direct propagation, reverse propagation, and gradient calculation steps. They are often parallelized in shared memory systems using OpenMP API and employing an OpenMP loop scheduler, as shown in Section 2.3. Aspects such as memory size, number of cores, and input size may influence the choice of an ideal scheduler.
To minimize FWI’s execution time, we propose using PATSMA with CSA to adjust the chunk size for OpenMP dynamic scheduling of parallel loops in an FWI. We use the time to solve the 3D wave equation as PATSMA’s cost function and the chunk size for OpenMP dynamic scheduling as its parameter.
Our FWI implementation is based on Assis et al. (2019). This version of FWI uses optimal checkpointing (Symes, 2010) with the library proposed by Griewank and Walther (2000). As optimal checkpointing also computes the solution of the wave equation, we also use the adjusted chunk size to schedule checkpointing loops dynamically.
Algorithm 1 illustrates an FWI with our proposed strategy. We propose executing PATSMA only on the first time step ($t_{i}==0$) of the first FWI shot (Line 10). PATSMA’s function $autotuning()$ (Line 11) optimizes the chunk size for loops of the wave propagator. The resulting chunk size is used in the forward propagation (Line 14), reverse propagation (Line 23), and in the re-computation of the forward field via checkpointing (Line 27). We chose those three sets of loops because they represent the most computationally expensive steps for an FWI and implement the wave equation solution, allowing the reuse of the chunk size obtained by PATSMA.
Algorithm 1 Full Wavefield Inversion with the proposed autotuning
1:Begin time measurement
2:Read the initial model $m_{0}$
3:Read the number of FWI iterations ($N_{\text{fwi}}$)
4:Read the number of time iterations ($N_{\text{s}}$)
5:Initialize other parameters for checkpointing, FWI, and PATSMA
6:for $k$ = 00 to $N_{\text{fwi}}$ do
7:     # Begin OpenMP parallel section
8:     for all shots do
9:         Read the observed data of the shot
10:         if it is the first shot and $k$ == 00 then
11:              chunk size = $autotuning()$
12:         end if
13:         for each time iteration $t_{i}={0\cdots N_{\text{s}}-1}$ do
14:              #OpenMP parallel loop directive using dynamic scheduler with the adjusted chunk size
15:              for all points in the domain of space do
16:                  Compute the direct wave field for the model $m_{k}$
17:              end for
18:              if $t_{i}$ it’s a checkpoint then
19:                  Save checkpoint
20:              end if
21:         end for
22:         for each time iteration $t_{i}={N_{\text{s}}-1}\cdots 0$ do
23:              #OpenMP parallel loop directive using dynamic scheduler with the adjusted chunk size
24:              for all points in the domain of space do
25:                  Calculate the adjunct wavefield for the model $m_{k}$
26:              end for
27:              Re-compute direct wavefield in $t_{i}$ from checkpoints using dynamic schedule with the adjusted chunk size
28:              #OpenMP parallel loop using static scheduler
29:              Calculate the objective function gradient
30:         end for
31:     end for
32:     Calculate search direction
33:     Determine the step size
34:     Update the model $m_{k+1}$
35:     # End OpenMP parallel section
36:end for
37:Final time measurement
The initial solution (chunk size) of each CSA optimizer is chosen randomly in the range [$50$, $N_{i}/N_{t}$]. We disregard chunk sizer inferior to $50$ due to the peak overhead generated to distribute them dynamically, as shown by Fernandes et al. (2018); Barros et al. (2018). Chunk sizes larger than the size of the standard static distribution chunk ($N_{i}/N_{t}$) are also not taken into account because they would make the number of chunks less than or equal to the number of threads, thus forcing it to be a static distribution.
For each iteration of the CSA, each optimizer measures only the execution time of the first time step of the forward propagation using its current chunk size. According to Assis et al. (2020), the execution time of the first time step can represent the total execution time of the propagation. This first step is executed twice, and only the elapsed time of the second repetition is recorded to avoid cache population effects. The CSA then uses those time measurements as the cost function values and generates the next set of solutions.
4 Results
This section details the tests we ran to measure the performance of the proposed strategy compared to standard OpenMP loop scheduling. Subsections 4.1 and 4.2 present the computational environments and the values of the parameters we used in our tests, respectively. In Subsection 4.3, we describe the tests we performed and present and discuss their results.
4.1 Computing Environments
We used different computational environments for the performance tests of the proposed autotuning, as described below:
  • OPT3 - VM.Optimized3.Flex is an Oracle virtual machine with an Intel Xeon 6354 processor with 18 cores, a base frequency of 3.0 GHz, a maximum turbo frequency of 3.6 GHz, and 256 GB of RAM.
  • SD - represents a node of SDumont’s supercomputer located at the National Center for Scientific Computing (LNCC). This node contains 2 Intel Xeon E5-2695v2 Ivy Bridge CPUs (12c @2.4GHz) totaling 24 cores and 64 GB of RAM.
  • NPAD - NPAD represents a node of the NPAD supercomputer located at the Federal University of Rio Grande do Norte (UFRN). This node contains 2 Intel Xeon CPU E5-2683 v4 @ 2.10GHz CPUs, for a total of 32 cores, and 512 GB of RAM
  • STD3 - VM.Standard3.Flex is an Oracle virtual machine with an Intel Xeon Platinum 8358 processor with 32 cores, a base frequency of 2.6 GHz, a maximum turbo frequency of 3.4 GHz, and 512 GB of RAM.
  • STDE4 - VM.Standard.E4.Flex is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 64 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, and 1024 GB of RAM.
  • DENSE - BM.DenseIO.E4.128 is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 128 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, and 2048 GB of RAM.
4.2 Test parameters
In terms of the FWI parameterization, for all tests conducted here, we used a peak frequency ($f_{peak}$) of $10Hz$, a time sampling interval of $1ms$, a total of $2458$ time steps, a spatial resolution of $\Delta x1=\Delta x2=\Delta x3=10m$, and an absorbing boundary thickness of $25$ grid points in all directions of the 3D mesh. We constructed $c(x)$ using a sphere with a Gaussian perturbation, where the upper layers have velocities of $2500m/s$ and its core has a maximum velocity of $3500m/s$. We used $3$ velocity models of dimensions $(n1,n2,n3)=(100,400,400)$, $(200,400,400)$, and $(400,400,400)$ where $n1$, $n2$, and $n3$ are the number of samples for the spatial dimensions $x1$, $x2$, and $x3$. The model with $n1=400$ is shown in Fig. 1. The others are slices taking the first $n1$ points of the model from Fig. 1 in the $x1$ dimension.
Figure 1: Velocity model with dimensions $n1=n2=n3=400$.
This paper uses CSA’s parameters discussed and tested by Assis et al. (2020). Table 1 contains the parameters used in our tests.
Table 1: CSA parameters where ${T_{0}^{gen}}$ and ${T_{0}^{ac}}$ are the initial generation and acceptance temperatures, respectively, $N$ is the total number of iterations, and $m$ is the number of optimizers.
${T_{0}^{gen}}$ ${T_{0}^{ac}}$ $N$ $m$
$100$ $0.9$ $40$ $4$
4.3 Performance analysis
The first set of tests aims to verify if it is possible to reuse the chunk of the first shot for the next ones. The model size chosen was $(n1,n2,n3)=(200,400,400)$ and we used $1$, $2$, $4$, $8$ and $16$ shots. These tests were performed on the NPAD machine, and each value presented represents a median of $5$ points.
Fig. 2 shows that, in all cases, the proposed strategy presented better performance, with speedups of up to $7.98\%$. Increasing the number of shots, we observe that the proposed autotuning performance remains stable because its overhead does not increase, being similar to the first shot. In this set of tests, the overhead of the proposed method remained below $1.2\%$ in all cases. The stable speedup indicates that the chunk size adjusted for the first shot can be reused in the following shots.
Figure 2: Speedup for FWI using the proposed autotuning compared to the OpenMP static and guided schedulers on the NPAD machine, with $1$, $2$, $4$, $8$, and $16$ shots. For the OpenMP schedulers, the chunk size was not explicitly set. The velocity model size for these tests was $(n1,n2,n3)=(200,400,400)$. Each point is a median of five executions.
In another set of tests, we measured the performance of the proposed strategy as we increased the problem size and computational resources. Fig. 3 illustrates the performance of the proposed autotuning compared to the standard OpenMP schedulers static and guided on the set of six computational environments mentioned in Subsection 4.1, varying $n1=\{100,200,400\}$. For the standard OpenMP schedulers, the chunk size was not explicitly set. Each point in the graph is a median of five samples. Fig. 3 shows that the proposed method had a superior performance in all scenarios. It also shows that, in most cases, the more cores a machine has, the higher the obtained speedup. The reason is that the higher the number of cores, the smaller the search space, implying fewer possible solutions (see Section 3). That way, CSA is more likely to find a near-optimal solution.
Figure 3: Single-shot FWI speedup using the proposed autotuning compared to OpenMP (a) static and (b) guided schedulers on $5$ machines (OPT3, SD, NPAD, STD3, STDE4, and DENSE), for three input sizes, $(n1,n2,n3)=(100,400,400)$, $(200,400,400)$, and $(400,400,400)$. For OpenMP schedulers, the chunk size was not explicitly set. Each point is a median of at least five runs.
As we can see, the speedups related to the static scheduler (Fig. 3a) are higher than those related to the guided scheduler (Fig. 3a). A poor performance of the static scheduler for large problem sizes can explain that phenomenon. The larger the problem size for the default static scheduler, the larger its chunk size (see Section 2.3). That means reduced data locality and, consequently, reduced cache reuse. Compared to the static scheduler, the FWI using the proposed strategy obtained a speedup of up to $70.46\%$.
Also, the wavefield is initially zero for each of the FWI’s wave propagations. As the wave propagation simulation runs, only a subset of the domain has non-zero wave pressure values. Because we use compilation optimization flags, the processor might run operations with zero values faster. When using the static scheduler, each thread is assigned a fixed block of iterations to process, regardless of the time it takes to execute each iteration. Since blocks’ load may vary, each block’s processing time may differ. If a thread is faster than the others, it can become idle while waiting for the other threads to finish their work. Therefore, the performance of the static scheduler is dominated by the slowest thread.
On the other hand, guided scheduler reduces threads’ idle time. Chunks are distributed dynamically, and by the end of the execution, chunk sizes are smaller. That way, the maximum idle time a thread may have is the time for processing the last chunk. guided scheduler can be more efficient than dynamic scheduler in cases where the ideal block size is unknown. However, the wave propagation effect can affect its performance. Should the first chunk (the largest) contain most of the non-zero wavefield, it will dictate the performance. In addition, guided scheduler has an overhead in managing the block size, which does not occur with dynamic scheduler. FWI using the proposed approach had a speedup of up to $15.27$% compared to using guided.
In our last set of experiments, we compared the runtime of an FWI using the proposed autotuning, the OpenMP static and guided scheduler with their default chunk size values, and the OpenMP dynamic scheduler with the chunk size defined by Mohammed et al. (2022):
$chunk=\left\lfloor\frac{N}{2^{f}\times 2P}\right\rfloor,$ (2)
where
$f=\left\lfloor log_{2}\left(\frac{N}{P}\right)\times{\frac{1}{\phi}}\right\rfloor,$ (3)
$N$ denotes the number of loop iterations, $P$ is the number of OpenMP threads, and $\phi=1.618$ as defined by Mohammed et al. (2022).
(a) $(n1,n2,n3)=(100,400,400)$
(b) $(n1,n2,n3)=(200,400,400)$
(c) $(n1,n2,n3)=(400,400,400)$
Figure 4: Single-shot FWI runtime using the proposed autotuning compared to the default OpenMP static and guided schedulers, and the OpenMP dynamic using the chunk size proposed by Mohammed et al. (2022). This set of experiments was performed on NPAD machine for three input sizes, $(n1,n2,n3)=$ (a) $(100,400,400)$, (b) $(200,400,400)$, and (c) $(400,400,400)$. Each point is a median of at least five runs.
As Fig. 4 shows, our proposed method presented the lowest runtimes in all scenarios tested in this set of experiments. In comparison to the case where the chunk size is the one defined by Mohammed et al. (2022), our autotuning performed about $4\%$ faster for the three input sizes we tested.
5 Related Work
Autotuning has been widely studied. Some authors present methods to adjust workflows automatically. HoseinyFarahabady et al. (2020) proposed an autotuning method to improve the data flow of stream processing applications based on continuously collected runtime information (e.g., available resources and incoming traffic rate). Shu et al. (2021) proposed the Component-based Ensemble Active Learning (CEAL). This autotuning method combines machine learning techniques with knowledge of the in-situ workflow structure to enable automated workflow configuration with limited performance measurements. Our proposal aims to autotune applications, unlike HoseinyFarahabady et al. (2020); Shu et al. (2021).
Application autotune has been exploited from several perspectives. Padoin et al. (2014, 2017) employ load balancing techniques to control processor frequency, improving the energy efficiency of parallel applications on multicore systems. Both works use processor frequency tuning techniques to slow down less loaded cores, saving energy. Rocha et al. (2020) presents AtTune, a heuristic-based framework to optimize the parallel execution of applications, tuning the number of processes/threads and CPU frequency levels. Bez (2021) proposes to use machine learning techniques to automatically detect I/O access patterns at runtime for applications in High-Performance Computing (HPC) platforms. Based on those patterns, it is possible to dynamically adjust I/O parameters and allocate resources more efficiently to improve the system’s overall performance. Bağbaba et al. (2021) presented an autotuning approach based on I/O monitoring and predictive modeling. This approach can find a set of I/O parameter values in a given system and application use case. This autotuning is applied to parallel I/O stacks in distributed systems using the MPI-IO ROMIO library. Robert et al. (2020) developed an optimization tool called SHAMAN, which provides a ready-to-use web application to perform black-box autotuning of custom computer components in a distributed system for a specific application executed through the Slurm workload manager. Different from Padoin et al. (2014, 2017); Rocha et al. (2020); Bez (2021); Bağbaba et al. (2021); Robert et al. (2020), our autotuning works on code-level optimization.
Several works proposed approaches for code-level optimization autotuning. Kale et al. (2014) proposed to improve scheduling by combining static and dynamic approaches in symmetric multiprocessing (SMP) machines, emphasizing spatial locality. Different from Kale et al. (2014), we propose an autotuning for any parallel application using OpenMP.
Katagiri et al. (2014, 2015) presented ppOpen-AT, a framework for directive-driven code optimization such as loop split and loop fusion. Menon et al. (2020) presented HiPerBOt, a configuration selection framework based on Bayesian optimization, to identify application and platform-level parameters that result in high-performance configurations. Liu et al. (2021) present GPTune, an autotuning based on multitask learning and Bayesian optimization suitable for tuning exascale application codes. Kimovski et al. (2021) developed a machine learning-based framework to optimize exascale applications. The AutoTuner can adjust the application’s parameters to improve performance when detecting an anomaly. Roy et al. (2021) presented Bliss, an algorithm that uses Bayesian optimization to find the near-optimal parameter configuration. Bliss must run the application a few times to determine the cost function. Milani (2020) proposed policies to choose the best among user-defined code versions automatically. In common, Katagiri et al. (2014, 2015); Menon et al. (2020); Liu et al. (2021); Kimovski et al. (2021); Roy et al. (2021); Milani (2020) need to run the application under a variety of different configurations to learn the correlation between an application’s parameters and its performance, which can be prohibitive for computationally expensive applications such as FWI. We propose using the time to compute the first iteration of the laplacian as an estimative of the time of the FWI.
Sakurai et al. (2020) proposed two autotuning functions for heterogeneous computing architectures. The first function may perform loop transformations to adjust the position of directives in OpenMP loops automatically. The second function dynamically changes the number of threads. Bak et al. (2018) introduced a load-balancing method based on integrating Charm++ and OpenMP to distribute user-created tasks dynamically. Kruse et al. (2020) addressed compiler performance optimization through a loop transformation autotuning. Their proposal is based on search trees to explore the loop transformation space and uses pragmas inserted in the source code to apply these transformations. Our proposal aims to optimize parallel loops scheduling, unlike Sakurai et al. (2020); Bak et al. (2018); Kruse et al. (2020).
Many authors proposed autotuning methods suitable for optimizing parallel loops scheduling. Dutta et al. (2022) presented preliminary work on using deep machine learning to identify loop patterns and adjust loop parameters with OpenMP. As Dutta et al. (2022) use machine learning, it is necessary to train it before execution. Wood et al. (2021) introduced Artemis, an online framework that dynamically tunes the execution of parallel regions by training optimizing models. Our proposed method requires no training, unlike Dutta et al. (2022); Wood et al. (2021).
Seiferth et al. (2020) introduced an approach that autotune applications by classifying optimization approaches based on a performance analytical model. Andreolli et al. (2014, 2015) brought an approach to automatically tune seismic applications by compiling and executing each set of parameters chosen by a genetic algorithm, including load block size and compilation flags. Seiferth et al. (2020); Andreolli et al. (2014, 2015) methods perform autotuning statically. That way, they may encounter a system state different from the runtime, as aspects such as memory availability, resource usage and access, workload, and cores’ load may vary over time. On the other hand, our method performs autotuning dynamically at runtime, allowing it to adjust the application for the system’s current state.
Dennis Booth and Allen Lane (2022) presented iCh, an autotuning method for OpenMP loop schedulers (guided and dynamic) that uses low-cost heuristics to reduce load imbalance. iCh uses on-demand work-stealing as its primary load-balancing mechanism and a distribution system with local queues for threads. By using work-stealing, iCh may have extra parallel overhead as it has to manage queues of tasks and thread synchronization. On the contrary, our proposal uses the OpenMP dynamic schedule.
Rasch et al. (2021) presented ATF (Automatic Tuning Framework), a general approach for automatically tuning programs with interdependent tuning parameters. It focuses on optimizing the efficiency of automatic tuning by providing optimized mechanisms for generating, storing, and exploring the search space of parameters. ATF observes past constraints to generate a search space, while in our proposal, the search field is defined in advance without the need for the generation step. ATF uses a multidimensional search strategy, while we use CSA, which, as a metaheuristic, seeks to reduce the search space.
Mohammed et al. (2022) introduced three methods for automated scheduler selection and an automatic chunk size adjustment for the OpenMP auto scheduler. They provide better performance by adapting to unpredictable variations in the application and system during execution. Mohammed et al. (2022) propose an equation to compute the chunk size based on the number of threads and the problem size. Doing so may disregard other essential factors, such as hardware architecture and software behavior (e.g., memory access pattern). Using PATSMA with CSA to optimize the chunk size, our proposal will more likely consider all relevant factors. In fact, for the experiments shown in Fig. 4, an FWI performed better using a chunk size obtained with our method rather than the chunk size proposed by Mohammed et al. (2022).
Assis et al. (2020) proposed using PATSMA with CSA to optimize the chunk size of parallel loops of a reverse time migration (RTM). Here, we expand Assis et al. (2020)’s approach to make it suitable for an FWI. FWI poses an extra challenge to autotuning methods as it has higher algorithmic complexity and a potentially different input for each iteration. This paper also increases reproducibility as tests were also performed in cloud instances. Finally, we present comparative tests with a state-of-the-art autotuning method.
6 Conclusion
This work proposed an autotuning strategy based on the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) to find an ideal chunk size for a dynamic OpenMP scheduler. This approach was designed to work in different computing environments, regardless of parameters such as number of threads, processors, and amount of memory. We applied the proposed mechanism to a Full Waveform Inversion (FWI) algorithm, reducing its execution time.
Experiments varying the input size and in different computing environments showed that the proposed autotuning outperforms the default OpenMP static and guided schedulers in all tested scenarios. For those experiments, the proposed strategy achieved speedups of up to $70.46\%$. Those results demonstrate that our method adapts well to different problem sizes and computing environments, outperforming the OpenMP standard schedulers.
Another set of experiments applied to an FWI showed the performance of the proposed method compared to the OpenMP static and guided schedulers for different numbers of seismic shots. In this set of tests, the proposed autotuning also outperformed the OpenMP schedulers in tested scenarios, reaching up to $7.98\%$ speedup. In this set of tests, it was observed that the speedup remained stable because the overhead did not increase with the increase in the number of shots. It is possible to conclude that the chunk size adjusted for the first shot can be reused for the following shots and iterations of FWI. In all cases, the overhead was below $1.2\%$. The results also show that the proposed method is scalable, presenting higher speedups for tests with larger input sizes and on architectures with more cores.
Finally, we have compared the execution times of an FWI using the OpenMP dynamic with the chunk size determined by our autotuning and by the formula proposed by Mohammed et al. (2022). For different input sizes, our method performed about $4\%$
7 Acknowledgments
The authors gratefully acknowledge support from Shell Brazil through the project “Novas metodologias computacionalmente escaláveis para sísmica 4D orientado ao alvo em reservatórios do pré-sal” at the Universidade Federal do Rio Grande do Norte (UFRN) and the strategic importance of the support given by ANP through the R&D levy regulation. This work was partly supported by Oracle Cloud credits and related resources the Oracle for Research program provided. The authors also acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) and the High-Performance Computing Center at UFRN (NPAD/UFRN) for providing HPC resources of the SDumont and NPAD supercomputers, which have contributed to the research results reported within this paper. Felipe Silva has been an undergraduate researcher through the PIVIC and PIVIC-Af programs at UFERSA.